OpenStudy (anonymous):

1. 7x + 2(8-x) = 51 2.(1/9)(0-x) + 2 = 19/9 3. (4/3)(x+6) - x= 8 4.(1+x)/3 = 2 5. x(3+1) -5x = 9

OpenStudy (anonymous):

1. 7x + 2(8-x)= 51 7x + 16-2x =51 7x - 2x = 51-16 5x = 35 x = 7 2. (1/9)(0-x) + 2 = 19/9 (times 9 throughout) 0-x +18 = 19 -x = 19-18 -x= 1 x = -1 3. (4/3)(x+6) - x= 8 (multiply 3 throughout) 4x + 24 - 3x = 24 x = 0 4. (1+x)/3 = 2 1+x = 6 x = 5 5. x(3+1) -5x = 9 3x+x - 5x = 9 -x = 9 x= -9